# Adaptive refinement for eigenvalue problems based on an associated source problem

### v1

@@ 5. The Algorithms
Algorithm 1 is standard, and will not be discussed in any detail here, apart from clarifying that, although it targets the $j$th eigenpair, the first $j$ eigenpairs are computed to ensure that we really have the $j$th in hand to drive the refinement. Both Algorithms 2 and 3 target the first $M$ eigenpairs of the operator, and both use the same criterion for convergence. They differ (significantly) in how they go about marking element for refinement and deciding how to refine them. In the case of Algorithm 3, these decisions are based on the current approximation of the landscape function $u_{h}$, as described in Algorithm 7. For Algorithm 2, all of the computed eigenpairs are used in some way to determine which elements are to be refined, and how, and two natural strategies are given in Algorithms 5 and 6. The two key parameters in Algorithms 4-7 are $r$, which is the percentage of element to be marked for refinement, and $\mathrm{tol}_{\mathrm{ana}}$, which dictates the choice of $h$ versus $p$ refinement for a marked element based on a computed measure of the local regularity of a function in that element. The default settings for these parameters for our experiments, are $r=10$% and $\mathrm{tol}_{\mathrm{ana}}=0.25$. The local regularity measure is computed using the method presented in [25], which is based on the coefficient decay of the local Legendre series expansion. This assessment takes place on line 4 of Algorithm 4, line 6 of Algorithm 5, line 5 of Algorithm 6, and line 4 of Algorithm 7.
@@ 6.2. Iterative solvers
In this section, we further reduce the CPU time by exploring different eigenvalue solvers. So far we used the ARPACK package [27] which is based on the Arnoldi method. ARPACK is very fast to compute multiple eigenpairs, however, in an adapting method, not all eigenpairs reach the desired tolerance on the same mesh. Therefore, it is common that on some meshes only a few of eigenpairs have not converged yet. ARPACK computes approximations of all eigenpairs, even of the ones already approximated well enough on previous meshes. In [20, 30], iterative eigenvalue solvers based on Picard’s and Newton’s methods are described. Even if these solvers might be slower than ARPACK in computing several eigenpairs, they can used to just compute a subset of the eigenpairs in the computed spectrum not necessarily contiguous in the spectrum.

### v2

@@ 5. The Algorithms
Algorithm 1 is standard, and will not be discussed in any detail here, apart from clarifying that, although it targets the $j$th eigenpair, the first $j$ eigenpairs are computed to ensure that we really have the $j$th in hand to drive the refinement. Both Algorithms 2 and 3 target the first $M$ eigenpairs of the operator, and both use the same criterion for convergence. They differ (significantly) in how they go about marking elements for refinement and deciding how to refine them. In the case of Algorithm 3, these decisions are based on the current approximation of the landscape function $u_{h}$, as described in Algorithm 7. For Algorithm 2, all of the computed eigenpairs are used in some way to determine which elements are to be refined, and how, and two natural strategies are given in Algorithms 5 and 6. The two key parameters in Algorithms 4-7 are $r$, which is the percentage of elements to be marked for refinement, and $\mathrm{tol}_{\mathrm{ana}}$, which dictates the choice of $h$ versus $p$ refinement for a marked element based on a computed measure of the local regularity of a function in that element. The default settings for these parameters for our experiments, are $r=10$% and $\mathrm{tol}_{\mathrm{ana}}=0.25$. The local regularity measure is computed using the method presented in [26], which is based on the coefficient decay of the local Legendre series expansion. This assessment takes place on line 4 of Algorithm 4, line 6 of Algorithm 5, line 5 of Algorithm 6, and line 4 of Algorithm 7.
@@ 6.2. Iterative solvers
In this section, we further reduce the CPU time by exploring different eigenvalue solvers. So far we used the ARPACK package [29] which is based on the Arnoldi method. ARPACK is very fast to compute multiple eigenpairs, however, in an adapting method, not all eigenpairs reach the desired tolerance on the same mesh. Therefore, it is common that on some meshes only a few of the eigenpairs have not converged yet. ARPACK computes approximations of all eigenpairs, even of the ones already approximated well enough on previous meshes. In [21, 32], iterative eigenvalue solvers based on Picard’s and Newton’s methods are described. Even if these solvers might be slower than ARPACK in computing several eigenpairs, they can be used to just compute a subset of the eigenpairs in the computed spectrum that are not necessarily contiguous in the spectrum.
@@ 8. Declarations
Competing Interests
The authors have no competing interests to declare that are relevant to the content of this article.
Data Availability
Data will be made available upon reasonable request.